OpenStudy (anonymous):

Reid and Maria both play soccer. This season, Reid scored 4 less than twice the number of goals that Maria scored. The difference in the number of goals they scored was 6. How many goals did each of them score?

OpenStudy (anonymous):

Let \(R\) and \(M\) denote the number of goals Reid and Maria scored, respectively. Then \[\begin{cases} R=2M-4\\ |R-M|=6 \end{cases}\] The second equation says either \(R-M=6\) or \(M-R=6\); we're not sure yet whether \(M\) or \(R\) is greater. Suppose it's \(R-M=6\), then \(R=M+6\) and subbing into the first equation of the system gives \[M+6=2M-4~~\iff~~M=10~~\Rightarrow~~R=16\] Suppose it's \(M-R=6\), then \(M=R+6\) and \[R=2(R+6)-4~~\iff~~\color{red}{R=-12}~~\Rightarrow~~\color{red}{M=-6}\] Clearly the first case must be true, since you can't score negative goals.